Kristin received a CD player for her birthday. She bought 3 CDs and received 5 others as gifts. What fraction of her total CDs did she buy and what fraction did she receive as gifts?

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Number of CDs Kristin bought from the market = 3
Number of CDs received as gifts = 5
Total number of CDs Kristin have = 3 + 5 = 8
∴Fraction of CD she bought = 3 / 8
∴Fraction of CDs received as gifts = 5 / 8

Thanks